# Swarming in the Wild: A Distributed Communication-less Lloyd-based Algorithm dealing with Uncertainties

### v1

@@ V Experimental Results
To validate our algorithm, we implemented it on real robotic platforms. We tested the proposed approach in two different scenarios: in the field with artificial obstacles, and in the forest. While in the former case we designed the distances between the obstacles and we relied on GNSS signal, in the forest we faced more challenging conditions both in terms of obstacle density, and in terms of localization, since we did not rely on GNSS signal. We started in a more controlled environment in order to have easier access to a ground truth measurement of the positions of the robots to evaluate the algorithm, and also for a higher repeatability of the experiment, because of the reduced complexity of the obstacle distribution.
Obstacle detection
@@
To detect the surrounding obstacles we relied on the RPLiDAR A3. In order to have a robust and safer perception system, we introduced a double layer of obstacles. The first layer of obstacles is given by the occupancy map built online by Hector [33], while the second layer is given by the raw pointcloud provided by the RPLiDAR A3. Given this information, we cluster the points on the basis of their spatial distribution. Then we compute the closest point to the robot position for each cluster, in order to estimate the position of the obstacles in the scene. By doing that, we exploit both the reactiveness of the raw data coming from the LiDAR pointcloud, and at the same time, the scan-matching based obstacle representation, which is provided by the Hector mapping algorithm.
Relative localization
The relative positions of the neighboring robots are obtained by relying on the UltraViolet Direction And Ranging (UVDAR) technology [34]. The sensor system operates by transmission of modulated ultraviolet light from active on-board markers. These markers consist of blinking ultraviolet LEDs and their signal is retrieved by on-board cameras with tailored filters. Each of these markers is individually identifiable according to its blinking signal, and retrieving a set of them corresponding to a specific Unmanned Aerial Vehicle (UAV) allows the system to estimate the UAV’s relative pose.
V-A Experiments in a field
We tested our approach on $N=3$ aerial robots in the field with artificial obstacles. We considered $\delta_{i}=1$ (m), $\delta^{o}=0.05$ (m), $r_{s,i}=6$ (m), $d_{1}=d_{3}=1$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=2.0$, $\Gamma_{ij}=12$ (m), $d_{u}=1.2$ (m), $\epsilon_{p}=\epsilon_{o}=2$, and the adjacency matrix that defines the flocking constraints is $A_{d}=\left[\begin{smallmatrix}0&0&1\\ 0&0&1\\ 1&1&0\end{smallmatrix}\right]$. In this first experiment, we positioned $N_{o}=7$ artificial obstacles that mimic a portion of the forest. The robots relied on GNSS. In accordance with theoretical findings, despite the presence of uncertainties in the perception, and tracking errors due to model mismatch and external disturbances, each robot safely converged to the goal regions maintaining the desired flocking constraints.
In Figure 6, we depict five snapshots at different time instants during the mission. The robots are depicted with colored circles depending on the time step, in transparency we depict $\delta_{i}+d_{u}$, which represents the amount of uncertainties that we are considering. The obstacles are depicted in black. The green crosses indicate the goal positions. As seen, the robots respected the flocking and the safety constraints and converged successfully to the goal regions.
Figure 6: Experimental result with $N=3$ aerial robots. We depicted $5$ snapshots in different time instants. Green crosses indicate the goal positions and the black circles indicate the obstacles. Transparent areas have radius equal to $d_{u}+\delta_{i}$.
Figure 7: On-board relative localization of UAV 1 and UAV 3 from the viewpoint of the UAV 2.
Figure 8: Distances measured by UVDAR system on UAV 2 from UAV 1 (blue) and UAV 3 (red). The flocking and safety constraints are denoted in magenta.
In Figures 7 and 8, we show the positions of the neighboring robots from the perspective of UAV $3$. The data provided by the onboard sensors is quite noisy, especially when distances between robots are significant and/or when occlusions due to the obstacles or other robots occur. Nevertheless, from Figure 8, it is clear that safety constraints are respected. On the other hand, the flocking constraint appears to exceed the limit of $\Gamma_{ij}=12$ (m) at certain times. By plotting the distances from the odometry data (Figure 9), it is clear that these are actually jumps in the measurements and not in the actual relative positions.
Figure 9: Distances between UAVs using odometry.
When this situation occurs, the uncertainty of the measure can be quite high, and since we set $d_{u}=1.2$ (m) as a constant value it may cause an appearance of constraint violation. From a practical point of view, the neighbor position is set to be at a maximum distance of $\Gamma_{ij}-d_{u}$, in order to maintain a well-defined cell, or we simply discard that value if the associated uncertainty is unreasonably large. Another solution would be to increase the value of $d_{u}$, but this would be at the cost of a drop of the performance. Similar evaluation was done for UAV $3$ (see Figure 1011). In this case the flocking constraints are active on both neighbors and, as it seen in Figure 9, the constraints are satisfied.
Figure 10: On-board relative localization of UAV 1 and UAV 2 from UAV 3 viewpoint.
Figure 11: Distances of UAV 1 (red) and UAV 2 (blue) measured by UVDAR system on UAV 3. The flocking and safety constraints are denoted in magenta.
V-B Experiments in a dense environment
@@
The second experimental deployment is conducted in the forest, where we do not have access to the GNSS signal, nor we can set a priori the obstacle positions. In this case, we change parameters to deal with more crowded conditions, according to the theoretical findings. The goal positions are set to be $40$ (m) ahead with respect the initial position of each robot, notice that we do not need to have a common reference frame, however we need to satisfy the requirement stated in Corollary 1, i.e., $\Gamma_{ij}>\|e_{i}-e_{j}\|>\Delta_{ij}$, $\forall i,j=1\dots N$ with $i\neq j$. We tested the algorithm with up to $N=4$, by selecting $\delta_{i}=0.3$ (m), $\delta^{o}=0.05$ (m), $r_{s,i}=6$ (m), $d_{1}=d_{3}=0.5$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=2.5$, $\Gamma_{ij}=12$ (m), $d_{u}=0.8$ (m), $\epsilon_{p}=\epsilon_{o}=1.052$, and the adjacency matrix that defines the flocking constraints is
$A_{d}=\begin{bmatrix}0&1&0&1\\ 1&0&1&0\\ 0&1&0&1\\ 1&0&1&0\end{bmatrix}.$
@@
In Figure 12, we show the trajectory of UAV $2$, the positions of the other robots measured by UVDAR system, and the map built by UAV $2$ by using Hector SLAM during the mission. In Figure 13, we depict the distances between UAV $2$ and the other robots according to the UVDAR measurements. Similar considerations as in the previous experiments apply also in this case. In Table III, we report the average velocities, the maximum velocities, and the time to reach the goal for each UAV during the mission. To verify the success of the mission, we report in Figure 15 the actual trajectories for all the 4 robots in the mission, while in Figure 14 we show the distances between the UAVs during the mission. As we expected the safety constrained and the flocking constraints (according to the adjancency matrix $A_{d}$) are satisfied.
@@
To show reliability and robustness to parameter changes, we report in Table IV the results obtained from other four experiments in the forest, with $N=3$ UAVs. Since the system is intrinsically chaotic it does not make sense to compare the results obtained, even if, as we can expect, a relation between $\beta^{D}$ and the time to reach the goal can be inspected from Table IV. Nevertheless, through the table we want to show the repeatability of the experiments and the robustness to parameter changes. The parameters for each experiment were set as follows: ##\##experiment $=[1,2,3,4]$, $N=[3,3,3,3]$, $\delta_{i}=[0.3,0.3,0.3,0.3]$ (m), $\delta^{o}=[0.05,0.05,0.05,0.05]$ (m) $r_{s,i}=[6,6,6,6]$ (m), $d_{1}=d_{3}=[0.5,0.5,0.5,0.5]$ (m), $d_{2}=d_{4}=[1,1,1,1]$ (m), $\beta_{i}^{D}=[2.0,2.0,1.0,0.6]$, $\Gamma_{ij}=[12,10,10,10]$ (m), $d_{u}=[0.75,0.75,0.75,0.75]$ (m), $\epsilon_{p}=\epsilon_{o}=[1.052,1.052,1.010,1.052]$, and the adjacency matrix that defines the flocking constraints is the same for all the four experiments, $A_{d}=\left[\begin{smallmatrix}0&0&1\\ 0&0&1\\ 1&1&0\end{smallmatrix}\right].$
@@
Figure 12: Experiment in the forest. The trajectory of UAV $2$ is depicted in green. The positions measured through UVDAR by UAV $2$ are depicted in orange (UAV $1$), blue (UAV $3$), and red (UAV $4$). The map is built by UAV $1$ by means of Hector mapping is also represented, in particular the free space is white, obstacles are black and the unexplored area is grey.
@@
Figure 14: Distances between UAVs using odometry.
Figure 15: Experiment in the forest. The trajectory of UAV $1$ is depicted in orange, UAV $2$ in green, UAV $3$ in blue, and UAV $4$ in red. The map built by UAV $2$ by means of Hector mapping is also represented, in particular the free space is white, obstacles are black and the unexplored area is grey.
@@ VI Conclusions
In this paper, we presented a novel distributed algorithm to swarm reliably and effectively in dense cluttered environments such as a forest, accounting for tracking errors and measurements uncertainties. We provide sufficient conditions for the theoretical guarantees of safety, proximity maintenance and convergence, compare it with state of the art algorithms, and finally we validate our solution through experiments in the field and in dense environment such as a forest. Our approach presents some imitations; the algorithm is limited to work on a plane, i.e., the reference in the altitude for each robot is constant and equal. To ensure the convergence towards the goal region, we assume that the obstacles are at a certain distance one each other (see Corollary 1). When these conditions are not met, providing to the algorithm an higher level re-planner can be beneficial to enhance performance. Future research directions will be focused on an effective 3D extension of the Lloyd algorithm and to increasing the agility of the swarm.

### v2

@@ V Experimental Results
To validate our algorithm, we implemented it on real robotic platforms. To detect the surrounding obstacles we relied on the RPLiDAR A3. In order to have a robust and safer perception system, we introduced a double layer of obstacles. The first layer of obstacles is given by the occupancy map built online by Hector [33], while the second layer is given by the raw pointcloud provided by the RPLiDAR A3. Given this information, we cluster the points on the basis of their spatial distribution. Then we compute the closest point to the robot position for each cluster, in order to estimate the position of the obstacles in the scene. By doing that, we exploit both the reactiveness of the raw data coming from the LiDAR pointcloud, and at the same time, the scan-matching based obstacle representation, which is provided by the Hector mapping algorithm. At the same time we do not need to know the obstacles’ size, since we are considering the closest point, we can account for a small value for $\delta^{o}$. The relative positions of the neighboring robots are obtained by relying on the UltraViolet Direction And Ranging (UVDAR) technology [28]. The sensor system operates by detection of blinking ultraviolet LEDs. The blinking signal is retrieved by on-board cameras with tailored filters. Each UAV is individually identifiable according to its blinking frequency. The pose of the UAVs is inferred through computer vision techniques.
@@
The experimental deployment is conducted in the forest, where we do not have access to the GNSS signal. The goal positions are set to be $40$ (m) ahead with respect to the initial position of each robot, notice that we do not need to have a common reference frame. However, we need to satisfy the requirement stated in Proposition 1, i.e., $\Gamma_{ij}>\|e_{i}-e_{j}\|>\Delta_{ij}$, $\forall i,j=1\dots N$ with $i\neq j$. We tested the algorithm with up to $N=4$, by selecting $\delta_{i}=0.3$ (m), $\delta^{o}=0.05$ (m), $r_{s,i}=6$ (m), $d_{1}=d_{3}=0.5$ (m), $d_{2}=d_{4}=1$ (m), $\beta_{i}^{D}=2.5$, $\Gamma_{ij}=12$ (m), $d_{u}=0.8$ (m), $\epsilon_{p}=\epsilon_{o}=1.052$, and the adjacency matrix that defines the flocking constraints is $A_{d}=\left[\begin{smallmatrix}0&1&0&1\\ 1&0&1&0\\ 0&1&0&1\\ 1&0&1&0\end{smallmatrix}\right]$.
@@
In Figure 8-(a) we show the trajectory of UAV $2$, the positions of the other robots measured by UVDAR system, and the map built by UAV $2$ by using Hector SLAM during the mission. In Figure 6, we depict the distances between UAV $2$ and the other robots according to the UVDAR measurements. In Table III, we report the average velocities, the maximum velocities, and the time to reach the goal for each UAV during the mission. To verify the success of the mission, we report in Figure 8-(b) the actual trajectories for all the 4 robots in the mission, while in Figure 7 we show the distances between the UAVs during the mission. As we expected the safety constraints and the flocking constraints (according to the adjancency matrix $A_{d}$) are satisfied.
@@
To show reliability and robustness to parameter changes, we report in Table IV the results obtained from other four experiments in the forest, with $N=3$ UAVs. Since the system is intrinsically chaotic it does not make sense to compare the results obtained, even if, as we can expect, a relation between $\beta^{D}$ and the time to reach the goal can be inspected from Table IV. Nevertheless, through the table we want to show the repeatability of the experiments and the robustness to parameter changes. The parameters for each experiment were set as follows: #\#experiment =[1,2,3,4]=[1,2,3,4], $N=[3,3,3,3]$, $\delta_{i}=[0.3,0.3,0.3,0.3]$ (m), $\delta^{o}=[0.05,0.05,0.05,0.05]$ (m) $r_{s,i}=[6,6,6,6]$ (m), $d_{1}=d_{3}=[0.5,0.5,0.5,0.5]$ (m), $d_{2}=d_{4}=[1,1,1,1]$ (m), $\beta_{i}^{D}=[2.0,2.0,1.0,0.6]$, $\Gamma_{ij}=[12,10,10,10]$ (m), $d_{u}=[0.75,0.75,0.75,0.75]$ (m), $\epsilon_{p}=\epsilon_{o}=[1.052,1.052,1.010,1.052]$, and the adjacency matrix that defines the flocking constraints is the same for all the four experiments, $A_{d}=\left[\begin{smallmatrix}0&0&1\\ 0&0&1\\ 1&1&0\end{smallmatrix}\right].$
@@
Figure 7: Distances between UAVs computed based on odometries transformed to a common reference frame obtained through alignment of maps of individual UAVs in a post-prossessing phase.
Refer to caption Refer to caption
(a) (b)
Figure 8: Experiment in the forest. In (a) the trajectory of UAV $2$ is depicted in green. The positions measured through UVDAR by UAV 22 are depicted in orange (UAV 11), blue (UAV 33), and red (UAV 44). The map built by UAV $2$ by means of Hector mapping is also represented, in particular the free space is white, obstacles are black and the unexplored area is grey. In (b) we depicted the actual trajectories obtained in post-process phase through transformation of odometries to a common reference frame. The map was built by UAV 11.
@@ VI Conclusions
In this paper, we presented a novel distributed algorithm to swarm reliably and effectively in dense cluttered environments represented by a forest, accounting for tracking errors and measurements uncertainties. We provide sufficient conditions for the theoretical guarantees of safety, proximity maintenance and convergence, compare it with state of the art algorithms, and finally we validate our solution through experiments in the field and in dense environment such as a forest. Our approach presents some limitations; the algorithm is limited to work on a plane, i.e., the reference in the altitude for each robot is constant and equal. To ensure the convergence towards the goal region, we assume that the individual obstacles are separated by a certain minimum distance (see Proposition 1, notice that these are just sufficient conditions and not necessary). Providing to the algorithm a higher level re-planner can be beneficial to enhance performance. Future research directions will be focused on an effective 3D extension of the Lloyd algorithm and to increase the agility of the swarm.